# Diffusion Models as Network Optimizers: Explorations and Analysis

### v3

@@ I Introduction
Network optimization problems arise from the widespread need to maximize or minimize specific objectives while utilizing limited network resources [1]. In the Internet of Things (IoT) network, these problems vary in form depending on the scenario, aiming for optimal resource allocation with highly dense deployed IoT devices [2, 3]. Common resources include communication capacity, storage, computational power, and energy. The optimization objectives typically involve maximizing transmission rates [4, 5], minimizing total delay [6] and energy consumption [7, 8], reducing execution costs [9, 10, 11], or enhancing quality of service (QoS) [12, 13]. The complexity of real-world environments and application requirements introduces challenging mathematical properties to these problems, such as convexity and non-convexity, linearity and non-linearity, and solution spaces that may be discrete, continuous, or mixed-integer. Additionally, these problems can involve both differentiable and non-differentiable components. These properties impact not only the objective functions but also the constraints, resulting in a complex feasible solution space that makes solving such problems particularly challenging.
@@
Existing network optimization methods primarily include numerical approaches based on optimization theory [5, 8, 10, 12, 13, 14, 15] and fitting algorithms based on machine learning [9, 11, 16, 17, 18, 19, 4, 20, 21, 22, 23], with some work exploring the use of deep learning to enhance numerical methods [24, 25, 26]. For problems without complex characteristics, it is often straightforward to apply classical algorithms. However, when dealing with more typical challenges—such as non-convex, mixed-integer, multi-objective, or Pareto optimization—custom algorithm design becomes necessary [5, 8, 10, 12, 13, 14, 15]. This process demands a strong understanding of optimization theory and can be labor-intensive in developing effective solutions, which makes IoT network designers burden. In the case of fitting algorithms, standard supervised learning [9, 11, 16, 17, 18] has limited capability to directly learn the mapping from input to optimal solutions, particularly when dealing with non-differentiable objective functions or constraints. Under such conditions, full convergence is nearly impossible, making it difficult to meet complex optimization demands. As a result, reinforcement learning (RL) [4, 20, 21, 22, 23], which interacts with the solution space and environment to make decisions, has become more commonly used. However, RL performance is heavily dependent on extensive interaction and exploration, which makes it challenging to train in high-dimensional, complex solution spaces of real IoT networks. Additionally, RL often suffers from issues such as reliance on reward function design and a tendency to get trapped in local optima. In numerical methods enhanced by deep learning [24, 25], certain components of traditional algorithms that are difficult to construct manually are replaced by learning-based approaches. While this can improve performance on some problems, it also demands more specialized designs compared to conventional numerical methods and hinders the efficiency of IoT construction. In summary, while conventional network optimization methods have achieved notable success, they often come with the drawbacks of heavy exclusive designs and inherent limitations in learning capacity.
The IoT network urges to address this issue for better performance. This paper proposes a novel approach focused on learning high-quality solution distributions for network optimization, which has the potential to overcome the above shortcomings. Accordingly, we introduce generative models for network optimization, especially diffusion models [27, 28], which have demonstrated strong performance in recent years.
@@
Figure 1: The learning results of generative models differ from those of discriminative models. Discriminative models focus on mapping $x$ directly to the optimal solution, but they often incur learning errors that are difficult to eliminate. In contrast, generative models learn the distribution of high-quality solutions for a given input $x$, with the optimal solution typically lying within this distribution.
@@ I-A Our Proposal
We first answer the intuitive question: “Why use generative models for network optimization?” Recall that the key difference between generative and discriminative models is that generative models learn the joint probability distribution $P(x,y)$, while discriminative models learn the conditional probability distribution $P(y|x)$. The core purpose of generative models is to capture a probability distribution and sample from it, whereas discriminative models focus on tasks like classification and regression. Additionally, generative diffusion models can leverage $P(y|x)$ to guide the sampling of specific types within the learned distribution [29, 30].
@@
However, we find that generative models possess a global awareness of the entire solution space for a given input, along with the ability to repeatedly sample solutions by learning the joint probability distribution from input to output, as illustrated in Fig. 1. Specifically, inspired by [31], we define the conditional generation target of the generative diffusion model as a probability distribution over solutions, $p_{\theta}(y|x)$, where $x$ is the given input and better solutions have higher probabilities. In this way, the model captures the feasible solution space corresponding to the input $x$ and probabilistically describes the quality of different solutions within it. Through repeated sampling, the model can approximate or even reach the global optimal solution, effectively closing the gap between traditional discriminative methods and the true optimal value. Notably, this approach to learning the solution distribution is insensitive to the complex characteristics of the original objective and constraints.
@@ II Literature Review and Analysis
This section introduces and compares the current works on network optimizers using diffusion models.
@@ II-B1 GDMs as Optimizers
In the networking domain, only a limited number of studies have directly applied GDMs as optimizers. [44] provides a case study using GDMs for a convex optimization problem, although it focuses more on using GDMs to enhance RL. [45] provides an early explanation of the mechanisms and advantages of using GDM as a network optimization solver, along with exploratory validation on several common network optimization problems. [46] employs discrete diffusion to generate graphs that minimize operational costs and utilizes continuous diffusion to create signals that evade monitoring. In [34], DIFUSCO’s approach is adapted to model Pareto optimization within Gantt chart representations, utilizing DDPM to generate Gantt charts for solution generation. However, as DIFUSCO argues, modeling problems as image representations may not enable explicit learning of the optimization problem itself.
In addition, GDM has shown significant potential in a class of problems known as black-box optimization [35, 36]. Black-box optimization primarily addresses problems with unknown objective functions, often arising from AI for Science (AI4S). These problems tend to be highly complex, meaning a model is typically tailored for a specific optimization setup. This one-shot nature of the GDM optimizer contrasts with the versatility and efficiency required in network optimization scenarios. Furthermore, recent work has provided an intriguing reverse perspective, applying optimization theory to the implementation of GDM conditional generation [49].
@@ II-B2 GDMs for RL Enhancement
In fact, the use of GDMs to enhance RL also originats from the AI field [47, 48], primarily in offline RL. These efforts primarily leverage GDMs’ strong ability to learn multi-modal data distributions—where “multi-modal” here refers to the probabilistic definition, describing complex data distributions for which a probability density function is nearly impossible to express analytically. This highlights the capacity of diffusion models to capture intricate distributions, underscoring the necessity and rationale for using generative models to learn distributions rather than traditional fitting algorithms.
In the network community, [37] discusses RL enhanced by GAI. [38] proposes D2SAC, which applies diffusion models as agents in RL frameworks for network optimization. Similarly, [39] discusses how diffusion models can streamline wireless network management. Their framework focuses on generating optimal network configurations, but many theoretical and experimental questions remain open. [40] introduces a diffusion-based approach for AI-Generated Network design (AIGN), automating intricate design tasks and optimizes network performance without requiring human expertise. [43] studies a multi-objective optimization problem of air communication security and energy efficiency using diffusion-enabled RL.
@@ II-B4 Other Generative Models as Optimizers
Notably, solving optimization problems with generative models suggests that a wide range of generative methods have the potential to function as optimizers, an area currently under active research. This includes approaches that leverage prompt engineering or fine-tuning of Large Language Models (LLMs) for solution optimization [59, 60, 61, 62, 63, 64]. Specifically, [59] first uses the trajectory of previous iterations combined with external prompts to guide an LLM toward the final solution of a target optimization problem. And [61] introduces LLMs into multi-objective optimization, proposing an LLM-Enabled Decomposition-based Multi-objective evolutionary Algorithm (LEDMA) to optimize Unmanned Aerial Vehicle (UAV) deployment and power control in integrated sensing and communication systems. Through careful prompt engineering to balance communication and sensing performance in non-convex problems, the model was guided to generate superior solutions iteratively. Similarly, [60] explores using LLM-driven evolutionary algorithms to solve the classic TSP problem. [62] applies prompt engineering to transform user voice requests into optimization problems, developing an LLM-based multi-agent framework for power scheduling in electric vehicle charging networks. They demonstrate how LLMs, with the help of prompts, can adjust scheduling strategies in real-time based on user preferences, enhancing energy efficiency while maintaining quality of service. [64] studies that LLMs can eliminate the dependency on prior knowledge and be seamlessly applied for various network management tasks. Additionally, [63] develops TelecomGPT, a specialized LLM framework for the telecommunications industry. By leveraging domain-specific datasets and fine-tuning prompts, TelecomGPT enhances its ability to tackle resource allocation and network management tasks, especially providing feasibility validation for translating natural language prompts into mathematical models.
@@ III-A Discussion of Generative Advantages
A typical network optimization problem can be formalized as an optimization problem: $\min_{y\in\mathcal{Y}}f(x,y),x\in\mathcal{X}$, where $x$ represents input parameters, $y$ denotes optimization variables, $f\rightarrow\mathbb{R}$ is the objective function, and $\mathcal{X}$ and $\mathcal{Y}$ are the domain and feasible region, respectively.
@@
First, we define the discriminative model $\phi$ and the generative model $\theta$. For a given $x$, their outputs are $h_{\phi}(x)$ and $y\sim p_{\theta}(y_{0}|x)$, respectively. Since model training cannot perfectly fit the target, we introduce a prediction error $e$. For simplicity and generality, we make no assumptions about the data format, dimensions, or magnitude of this error.
@@ III-A1 Lower Bound of the Expected Objective Function Value of Discriminative Model
The above definition means that $h_{\phi}(x)$ will deviate from the optimal solution $y^{*}$ and the degree is determined by $e$
$h_{\phi}(x)\in\partial{\mathrm{U}}(y^{*},e),$ (1)
where $\partial{\mathrm{U}}(y^{*},e)$ represents the boundary of the $e$ neighborhood of $y^{*}$.
Assume there is a probability $p$ that a solution outside $\mathrm{U}(y^{*},e)$ exceeds the objective function value of the optimal solution $f(x,y^{*})$ by a factor of $\sigma>1$
$\mathrm{Pr}\left(f(x,y)>f(x,y^{*})\sigma\ \mathrm{and}\ y\notin\mathrm{U}(y^{*},e)\right)=p,$ (2)
Now, we can derive the lower bound of the expected objective function value for the solution predicted by $\phi$
$\mathbb{E}\left[f(x,h_{\phi}(x))\right]\geq pf(x,y^{*})\sigma+(1-p)f(x,y^{*}),$ (3)
@@ III-A2 Lower Bound of the Expected Objective Function Value of Generative Model
Intuitively, we consider that the prediction error implies that the predicted optimal solution with the highest probability in the high-quality solution distribution generated by the model will deviate from $y^{*}$ by an amount $e$. However, although the predicted solution distribution of the generative model is shifted, it generally still retains a certain sampling probability for the true optimal solution. Therefore, we denote $p_{i}$ as the probability that a single sampled solution belongs to $\mathrm{U}(y^{*},e)$
$\mathrm{Pr}\left(y\sim p_{\theta}(y_{0}|x)\ \mathrm{and}\ y\in\mathrm{U}(y^{*},e)\right)=p_{i},$ (4)
Similarly, we denote $p_{o}$ as the probability that a single sampled solution is out of $\mathrm{U}(y^{*},e)$
$\mathrm{Pr}\left(y\sim p_{\theta}(y_{0}|x)\ \mathrm{and}\ y\notin\mathrm{U}(y^{*},e)\right)=p_{o},$ (5)
@@
$\mathbb{E}$ $\left[f(x,y\sim p_{\theta}(y_{0}|x))\right]\geq p_{i}f(x,y^{*})$ (6)
$\quad+p_{o}\left[pf(x,y^{*})\sigma+(1-p)f(x,y^{*})\right],$
@@ III-B Optimizer Formulation
Based on the above definition, we explain the noising and denoising processes of the discrete-time DDPM adopted in our work to learn the distribution of high-quality solutions $p(y_{0}|x)$, where the index 0 represents the original, uncorrupted data. Each pair $(x,y)$, representing an input and its corresponding optimal solution, forms our training dataset $\mathcal{D}$. Since $x$ has no influence during the noising process, we simplify this to $p(y_{0})$; it only serves as a fixed condition to guide the input during the denoising process.
For the noising process, the training data is gradually corrupted by Gaussian noise with mean and variance determined by the noise factor $(\alpha_{t})^{T}_{t=1}$. Meanwhile, $(\alpha_{t})^{T}_{t=1}$ are also regarded as variance scheduling factors in the diffusion steps. Given an uncorrupted training sample $y_{0}\sim p(y_{0})$, the noisy samples $y_{1},y_{2}...,y_{T}$ can be obtained through the following Markov process:
$p(y_{t}|y_{t-1})=\mathcal{N}(y_{t};\sqrt{\alpha_{t}}\cdot y_{t-1},(1-\alpha_{t})\cdot\boldsymbol{I}),\ \forall t\in{1,...,T},$ (8)
where $T$ represents the number of diffusion steps. Additionally, $\boldsymbol{I}$ stands for the identity matrix with the same dimension as $y_{0}$.
However, acquiring corrupted data $y_{t}$ for any step by totally depending on Eq. 8 necessitates an excessive number of noising iterations. Consequently, leveraging the recursive nature of this formula allows for the direct sampling of $y_{t}$ using the simplified approach outlined as
$p\left(y_{t}\mid y_{0}\right)=\mathcal{N}\left(y_{t};\sqrt{\bar{\alpha}_{t}}\cdot y_{0},\left(1-\bar{\alpha}_{t}\right)\cdot\boldsymbol{I}\right),$ (9)
where $t$ is drawn from a uniform distribution $\mathcal{U}(\{1,...,T\})$, and $\bar{\alpha}_{t}$ is defined as the product of ${\alpha}_{i}$ for $i$ ranging from $1$ to $t$. Consequently, Eq. 10 facilitates the sampling and learning of any noisy version ${y}_{t}$ directly.
Specifically, for any given sample $(x,y)$ in $\mathcal{D}$, we introduce noise $\epsilon_{t}$ sampled from a standard normal distribution $\mathcal{N}(0,\boldsymbol{I})$. And then randomly choose a time step $t$ from the uniform distribution and employ Eq. 10 to corrupt $y$ as
$y_{t}=\sqrt{\bar{\alpha}_{t}}y_{0}+\sqrt{\left(1-\bar{\alpha}_{t}\right)}\epsilon_{t},$ (10)
where the added noise $\epsilon_{t}$ is the prediction target of the model in denoising.
@@
$p_{\theta}\left(y_{t-1}\mid y_{t}\right)=\mathcal{N}\left(y_{t-1};\mu_{\theta}\left(y_{t},x,t\right),\Sigma_{\theta}\left(y_{t},x,t\right)\right),$ (11)
To handle conditional guidance from $x$ during the denoising process, we use the Classifier-Free Guidance mechanism [30] that jointly trains both conditional and unconditional models. Classifier-Free Guidance is able to handle some key challenges, such as the model’s neglect of conditional factors and the issues stemming from imprecise score estimation and classifier gradients highlighted in [29]. This approach does not require training any additional models; it simply controls the proportion of conditional samples during DDPM training and, in the denoising process, combines the conditional and unconditional noise predictions through weighted summation. These modifications to the classic DDPM training and sampling process involve only a few lines of code, making it a low-complexity, high-quality solution for conditional generation.
Specifically, in the training phase, a hyperparameter $p_{uncond}$ is introduced to determine to what extent the model should be trained without any condition (i.e., setting the condition input to $\varnothing$). This flexibility allows the model to learn robust representations in both conditioned and unconditioned contexts. In the sampling process, the intensity of conditional guidance is regulated by the weight $w$, which helps us achieve a refined score estimation. The resulting prediction noise is as
$\tilde{\epsilon}_{\theta}(y_{t},x,t)=(1+\omega)\epsilon_{\theta}(y_{t},x,t)-\omega\epsilon_{\theta}(y_{t},t),$ (12)
The approach in Eq. 12 adeptly mitigates the issue of underestimating conditions while requiring only a single model for training. Classifier-Free Guidance strikes a balance between maintaining diversity and adhering to conditions. Although some diversity may be sacrificed [65], this trade-off is advantageous in the context of optimization, facilitating convergence toward high-quality solution distributions. Ultimately, we utilize the noise prediction from Eq. 12 to guide the generation process in optimization inference, resulting in target solutions. Using the predicted noise $\tilde{\epsilon}_{\theta}$, the reconstruction function is defined as
$y_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(y_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\tilde{\epsilon}_{\theta})+\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\epsilon,\ \epsilon\sim\mathcal{N}(0,\boldsymbol{I}).$ (13)
Under the guidance of Eq. 13, the GDM can repeatedly sample from the target high-quality solution distribution $p(y_{0}|x)$ based on the input $x$, effectively approximating or reaching the optimal solution, as shown in Fig. 2.

### v5

@@ I Introduction
Network optimization problems arise from the widespread need to maximize or minimize specific objectives while utilizing limited network resources [1]. In the Internet of Things (IoT) network, these problems vary in form depending on the scenario, aiming for optimal resource allocation with highly dense deployed IoT devices [2, 3]. Common resources include communication capacity, storage, computational power, and energy. The optimization objectives typically involve maximizing transmission rates [4], minimizing total delay [5] and energy consumption [6, 7], reducing execution costs [8, 9, 10], or enhancing quality of service (QoS) [11, 12]. The complexity of real-world environments and application requirements introduces challenging mathematical properties to these problems, such as convexity and non-convexity, linearity and non-linearity, and solution spaces that may be discrete, continuous, or mixed-integer. Additionally, these problems can involve both differentiable and non-differentiable components. These properties impact not only the objective functions but also the constraints, resulting in a complex feasible solution space that makes solving such problems particularly challenging.
@@
Existing network optimization methods primarily include numerical approaches based on optimization theory [4, 7, 9, 11, 12, 13, 14] and fitting algorithms based on machine learning [8, 10, 15, 16, 17, 18, 19, 20, 21, 22], with some work exploring the use of deep learning to enhance numerical methods [23, 24, 25]. For problems without complex characteristics, it is often straightforward to apply classical algorithms. However, when dealing with more typical challenges—such as non-convex, mixed-integer, multi-objective, or Pareto optimization—custom algorithm design becomes necessary [4, 7, 9, 11, 12, 13, 14]. This process demands a strong understanding of optimization theory and can be labor-intensive in developing effective solutions, which makes IoT network designers burden. In the case of fitting algorithms, standard supervised learning [8, 10, 15, 16, 17] has limited capability to directly learn the mapping from input to optimal solutions, particularly when dealing with non-differentiable objective functions or constraints. Under such conditions, full convergence is nearly impossible, making it difficult to meet complex optimization demands. As a result, reinforcement learning (RL) [19, 20, 21, 22], which interacts with the solution space and environment to make decisions, has become more commonly used. However, RL performance is heavily dependent on extensive interaction and exploration, which makes it challenging to train in high-dimensional, complex solution spaces of real IoT networks. Additionally, RL often suffers from issues such as reliance on reward function design and a tendency to get trapped in local optima. In numerical methods enhanced by deep learning [23, 24], certain components of traditional algorithms that are difficult to construct manually are replaced by learning-based approaches. While this can improve performance on some problems, it also demands more specialized designs compared to conventional numerical methods and hinders the efficiency of IoT construction. In summary, while conventional network optimization methods have achieved notable success, they often come with the drawbacks of heavy exclusive designs and inherent limitations in learning capacity. In this paper, we propose a novel approach focused on learning high-quality solution distributions for network optimization, which has the potential to overcome the above shortcomings. Accordingly, we introduce generative models for network optimization, especially diffusion models [26, 27], which have demonstrated strong performance in recent years.
@@
Figure 1: The learning results of generative models differ from those of discriminative models. Discriminative models focus on mapping $\mathbf{x}$ directly to the optimal solution, but they often incur learning errors that are difficult to eliminate. In contrast, generative models learn the distribution of high-quality solutions for a given input $\mathbf{x}$, with the optimal solution typically lying within this distribution.
@@ I-A Our Proposal
We first answer the intuitive question: “Why use generative models for network optimization?” Recall that the key difference between generative and discriminative models is that generative models learn the joint probability distribution $P(\mathbf{x},\mathbf{y})$, while discriminative models learn the conditional probability distribution $P(\mathbf{y}|\mathbf{x})$. The core purpose of generative models is to capture a probability distribution and sample from it, whereas discriminative models focus on tasks like classification and regression. Additionally, generative diffusion models can leverage $P(\mathbf{y}|\mathbf{x})$ to guide the sampling of specific types within the learned distribution [28, 29].
@@
However, we find that generative models possess a global awareness of the entire solution space for a given input, along with the ability to repeatedly sample solutions by learning the joint probability distribution from input to output, as illustrated in Fig. 1. Specifically, inspired by [30], we define the conditional generation target of the generative diffusion model as a probability distribution over solutions, $p_{\theta}(\mathbf{y}|\mathbf{x})$, where $\mathbf{x}$ is the given input and better solutions have higher probabilities. In this way, the model captures the feasible solution space corresponding to the input $\mathbf{x}$ and probabilistically describes the quality of different solutions within it. Through repeated sampling, the model can approximate or even reach the global optimal solution, effectively closing the gap between traditional discriminative methods and the true optimal value. Notably, this approach to learning the solution distribution is insensitive to the complex characteristics of the original objective and constraints.
@@ II-B1 GDMs as Optimizers
In the networking domain, only a limited number of studies have directly applied GDMs as optimizers. [43] provides a case study using GDMs for a convex optimization problem, although it focuses more on using GDMs to enhance RL. [44] provides an early explanation of the mechanisms and advantages of using GDM as a network optimization solver, along with exploratory validation on several common network optimization problems. [45] employs discrete diffusion to generate graphs that minimize operational costs and utilizes continuous diffusion to create signals that evade monitoring. In [33], DIFUSCO’s approach is adapted to model Pareto optimization within Gantt chart representations, utilizing DDPM to generate Gantt charts for solution generation. However, as DIFUSCO argues, modeling problems as image representations may not enable explicit learning of the optimization problem itself. [58] employs graph diffusion to solve the optimization problems in Multi-access Edge Computing (MEC) networks and explores the convergence of the denoising solution generation process, but the method itself still has limited generality problems.
In addition, GDM has shown significant potential in a class of problems known as closed-box optimization [34, 35]. Closed-box optimization primarily addresses problems with unknown objective functions, often arising from AI for Science (AI4S). These problems tend to be highly complex, meaning a model is typically tailored for a specific optimization setup. This one-shot nature of the GDM optimizer contrasts with the versatility and efficiency required in network optimization scenarios. Furthermore, recent work has provided an intriguing reverse perspective, applying optimization theory to the implementation of GDM conditional generation [48].
@@ II-B2 GDMs for RL Enhancement
In fact, the use of GDMs to enhance RL also originates from the AI field [46, 47], primarily in offline RL. These efforts primarily leverage GDMs’ strong ability to learn multi-modal data distributions—where “multi-modal” here refers to the probabilistic definition, describing complex data distributions for which a probability density function is nearly impossible to express analytically. This highlights the capacity of diffusion models to capture intricate distributions, underscoring the necessity and rationale for using generative models to learn distributions rather than traditional fitting algorithms.
In the network community, [36] discusses RL enhanced by GAI. [37] proposes D2SAC, which applies diffusion models as agents in RL frameworks for network optimization. Similarly, [38] discusses how diffusion models can streamline wireless network management. Their framework focuses on generating optimal network configurations, but many theoretical and experimental questions remain open. [39] introduces a diffusion-based approach for AI-generated network design (AIGN), automating intricate design tasks and optimizing network performance without requiring human expertise. [42] studies a multi-objective optimization problem of air communication security and energy efficiency using diffusion-enabled RL.
@@ II-B4 Other Generative Models as Optimizers
Notably, solving optimization problems with generative models suggests that a wide range of generative methods have the potential to function as optimizers, an area currently under active research. This includes approaches that leverage prompt engineering or fine-tuning of Large Language Models (LLMs) for solution optimization [59, 60, 61, 62, 63, 64]. Specifically, [59] first uses the trajectory of previous iterations combined with external prompts to guide an LLM toward the final solution of a target optimization problem. And [61] introduces LLMs into multi-objective optimization, proposing an LLM-Enabled Decomposition-based Multi-objective evolutionary Algorithm (LEDMA) to optimize Uncrewed Aerial Vehicle (UAV) deployment and power control in integrated sensing and communication systems. Through careful prompt engineering to balance communication and sensing performance in non-convex problems, the model was guided to generate superior solutions iteratively. Similarly, [60] explores using LLM-driven evolutionary algorithms to solve the classic TSP problem. [62] applies prompt engineering to transform user voice requests into optimization problems, developing an LLM-based multi-agent framework for power scheduling in electric vehicle charging networks. They demonstrate how LLMs, with the help of prompts, can adjust scheduling strategies in real-time based on user preferences, enhancing energy efficiency while maintaining quality of service. [64] studies that LLMs can eliminate the dependency on prior knowledge and be seamlessly applied for various network management tasks. Additionally, [63] develops TelecomGPT, a specialized LLM framework for the telecommunications industry. By leveraging domain-specific datasets and fine-tuning prompts, TelecomGPT enhances its ability to tackle resource allocation and network management tasks, especially providing feasibility validation for translating natural language prompts into mathematical models.
@@ III-A Discussion of Generative Advantages
A typical network optimization problem can be formalized as an optimization problem: $\min_{\mathbf{y}\in\mathcal{Y}}f(\mathbf{x},\mathbf{y}),\mathbf{x}\in\mathcal{X}$, where $\mathbf{x}$ represents input parameters, $\mathbf{y}$ denotes optimization variables, $f\rightarrow\mathbb{R}$ is the objective function, and $\mathcal{X}$ and $\mathcal{Y}$ are the domain and feasible region, respectively.
@@
First, we define the discriminative model $\varphi$ and the generative model $\theta$. For a given $\mathbf{x}$, their outputs are $h_{\varphi}(\mathbf{x})$ and $\mathbf{y}\sim p_{\theta}(\mathbf{y}_{0}|\mathbf{x})$, respectively. Since model training cannot perfectly fit the target, we introduce a prediction error $\mathbf{e}$. For simplicity and generality, we make no assumptions about the data format, dimensions, or magnitude of this error.
@@ III-A1 Lower Bound of the Expected Objective Function Value of Discriminative Model
The above definition means that $h_{\varphi}(\mathbf{x})$ will deviate from the optimal solution $\mathbf{y}^{*}$ and the degree is determined by $\mathbf{e}$
$h_{\varphi}(\mathbf{x})\in\partial{\mathrm{U}}(\mathbf{y}^{*},\mathbf{e}),$ (1)
where $\partial{\mathrm{U}}(\mathbf{y}^{*},\mathbf{e})$ represents the boundary of the $\mathbf{e}$ neighborhood of $\mathbf{y}^{*}$.
Assume there is a probability $p$ that a solution outside $\mathrm{U}(\mathbf{y}^{*},\mathbf{e})$ exceeds the objective function value of the optimal solution $f(\mathbf{x},\mathbf{y}^{*})$ by a factor of $\sigma>1$
$\mathrm{Pr}\left(f(\mathbf{x},\mathbf{y})>f(\mathbf{x},\mathbf{y}^{*})\sigma\ \mathrm{and}\ \mathbf{y}\notin\mathrm{U}(\mathbf{y}^{*},\mathbf{e})\right)=p,$ (2)
Now, we can derive the lower bound of the expected objective function value for the solution predicted by $\varphi$
$\mathbb{E}\left[f(\mathbf{x},h_{\varphi}(\mathbf{x}))\right]\geq pf(\mathbf{x},\mathbf{y}^{*})\sigma+(1-p)f(\mathbf{x},\mathbf{y}^{*}),$ (3)
@@ III-A2 Lower Bound of the Expected Objective Function Value of Generative Model
Intuitively, we consider that the prediction error implies that the predicted optimal solution with the highest probability in the high-quality solution distribution generated by the model will deviate from $\mathbf{y}^{*}$ by an amount $\mathbf{e}$. However, although the predicted solution distribution of the generative model is shifted, it generally still retains a certain sampling probability for the true optimal solution. Therefore, we denote $p_{i}$ as the probability that a single sampled solution belongs to $\mathrm{U}(\mathbf{y}^{*},\mathbf{e})$
$\mathrm{Pr}\left(\mathbf{y}\sim p_{\theta}(\mathbf{y}_{0}|\mathbf{x})\ \mathrm{and}\ \mathbf{y}\in\mathrm{U}(\mathbf{y}^{*},\mathbf{e})\right)=p_{i},$ (4)
Similarly, we denote $p_{o}$ as the probability that a single sampled solution is out of $\mathrm{U}(\mathbf{y}^{*},\mathbf{e})$
$\mathrm{Pr}\left(\mathbf{y}\sim p_{\theta}(\mathbf{y}_{0}|\mathbf{x})\ \mathrm{and}\ \mathbf{y}\notin\mathrm{U}(\mathbf{y}^{*},\mathbf{e})\right)=p_{o},$ (5)
@@
$\mathbb{E}$ $\left[f(\mathbf{x},\mathbf{y}\sim p_{\theta}(\mathbf{y}_{0}|\mathbf{x}))\right]\geq p_{i}f(\mathbf{x},\mathbf{y}^{*})$ (6)
$\quad+p_{o}\left[pf(\mathbf{x},\mathbf{y}^{*})\sigma+(1-p)f(\mathbf{x},\mathbf{y}^{*})\right],$
@@ III-B Optimizer Formulation
Based on the above definition, we explain the noising and denoising processes of the discrete-time DDPM adopted in our work to learn the distribution of high-quality solutions $p(\mathbf{y}_{0}|\mathbf{x})$, where the index 0 represents the original, uncorrupted data. Each pair $(\mathbf{x},\mathbf{y})$, representing an input and its corresponding optimal solution, forms our training dataset $\mathcal{D}$. Since $\mathbf{x}$ has no influence during the noising process, we simplify this to $p(\mathbf{y}_{0})$; it only serves as a fixed condition to guide the input during the denoising process.
For the noising process, the original data is gradually corrupted by Gaussian noise with mean and variance controlled by the noise factor $(\alpha_{t})^{T}_{t=1}$. Given an uncorrupted training sample $\mathbf{y}_{0}\sim p(\mathbf{y}_{0})$, the noisy samples $\mathbf{y}_{1},\mathbf{y}_{2}...,\mathbf{y}_{T}$ can be obtained through the following Markov process:
$p(\mathbf{y}_{t}|\mathbf{y}_{t-1})=\mathcal{N}(\mathbf{y}_{t};\sqrt{\alpha_{t}}\cdot\mathbf{y}_{t-1},(1-\alpha_{t})\cdot\mathbf{I}),\ \forall t\in{1,...,T},$ (8)
where $T$ represents the number of diffusion steps. Additionally, $\mathbf{I}$ stands for the identity matrix with the same dimension as $\mathbf{y}_{0}$.
However, acquiring corrupted data $\mathbf{y}_{t}$ for any step by totally depending on Eq. 8 necessitates an excessive number of noising iterations. Consequently, leveraging the recursive nature of this formula allows for the direct sampling of $\mathbf{y}_{t}$ using the simplified approach outlined as
$p\left(\mathbf{y}_{t}\mid\mathbf{y}_{0}\right)=\mathcal{N}\left(\mathbf{y}_{t};\sqrt{\bar{\alpha}_{t}}\cdot\mathbf{y}_{0},\left(1-\bar{\alpha}_{t}\right)\cdot\mathbf{I}\right),$ (9)
where $t$ is drawn from a uniform distribution $\mathcal{U}(\{1,...,T\})$, and $\bar{\alpha}_{t}$ is defined as the product of ${\alpha}_{i}$ for $i$ ranging from $1$ to $t$. Consequently, Eq. 10 facilitates the sampling and learning of any noisy version ${\mathbf{y}}_{t}$ directly.
Specifically, for any given sample $(\mathbf{x},\mathbf{y})$ in $\mathcal{D}$, we introduce noise $\boldsymbol{\epsilon}_{t}$ sampled from a standard normal distribution $\mathcal{N}(0,\mathbf{I})$. And then randomly choose a time step $t$ from the uniform distribution and employ Eq. 10 to corrupt $\mathbf{y}$ as
$\mathbf{y}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{y}_{0}+\sqrt{\left(1-\bar{\alpha}_{t}\right)}\boldsymbol{\epsilon}_{t},$ (10)
where the added noise $\boldsymbol{\epsilon}_{t}$ is the prediction target of the model in denoising.
@@
$p_{\theta}\left(\mathbf{y}_{t-1}\mid\mathbf{y}_{t}\right)=\mathcal{N}\left(\mathbf{y}_{t-1};\mu_{\theta}\left(\mathbf{y}_{t},\mathbf{x},t\right),\Sigma_{\theta}\left(\mathbf{y}_{t},\mathbf{x},t\right)\right),$ (11)
To handle conditional guidance from $\mathbf{x}$ during the denoising process, we use the Classifier-Free Guidance mechanism [29] that jointly trains both conditional and unconditional models. Classifier-Free Guidance is able to handle some key challenges, such as the model’s neglect of conditional factors and the issues stemming from imprecise score estimation and classifier gradients highlighted in [28]. This approach does not require training any additional models; it simply controls the proportion of conditional samples during DDPM training and, in the denoising process, combines the conditional and unconditional noise predictions through weighted summation. These modifications to the classic DDPM training and sampling process involve only a few lines of code, making it a low-complexity, high-quality solution for conditional generation.
Specifically, in the training phase, a hyperparameter $p_{uncond}$ is introduced to determine to what extent the model should be trained without any condition (i.e., setting the condition input to $\varnothing$). This flexibility allows the model to learn robust representations in both conditioned and unconditioned contexts. In the sampling process, the intensity of conditional guidance is regulated by the weight $\omega$, which helps us achieve a refined score estimation. The resulting prediction noise is as
$\tilde{\boldsymbol{\epsilon}}_{\theta}(\mathbf{y}_{t},\mathbf{x},t)=(1+\omega)\boldsymbol{\epsilon}_{\theta}(\mathbf{y}_{t},\mathbf{x},t)-\omega\boldsymbol{\epsilon}_{\theta}(\mathbf{y}_{t},t).$ (12)
The approach in Eq. 12 adeptly mitigates the issue of underestimating conditions while requiring only a single model for training. Classifier-Free Guidance strikes a balance between maintaining diversity and adhering to conditions. Although some diversity may be sacrificed, this trade-off is advantageous in the context of optimization, facilitating convergence toward high-quality solution distributions. Ultimately, we utilize the noise prediction from Eq. 12 to guide the generation process in optimization inference, resulting in target solutions. Using the predicted noise $\tilde{\boldsymbol{\epsilon}}_{\theta}$, the reconstruction function is defined as
$\mathbf{y}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{y}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\tilde{\boldsymbol{\epsilon}}_{\theta})+\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},\ \boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I}).$ (13)
Under the guidance of Eq. 13, the GDM can repeatedly sample from the target high-quality solution distribution $p(\mathbf{y}_{0}|\mathbf{x})$ based on the input $\mathbf{x}$, effectively approximating or reaching the optimal solution, as shown in Fig. 2.